OpenStudy (anonymous):

list all possible rational zeros given by rational zeros therorem. do not check to see which actually are zeros. p(x)0 =2x^5+18x^3+8x^2-10

OpenStudy (anonymous):

ho ho ho

OpenStudy (anonymous):

fraction where numerator divides 10 and denominator divides 2

OpenStudy (amistre64):

the rational thrm zero thing is just taking all the numbers that can be made from factoring the first and last coeffs

OpenStudy (anonymous):

take out the o after p(x) its a typo

OpenStudy (amistre64):

factors of 10 ---------- factors of 2

OpenStudy (anonymous):

how would i write showing work

OpenStudy (amistre64):

youd have to factor out 10 and 2 and determine which numbers you can make ...out of them

OpenStudy (amistre64):

1,2,5,10 +- -------- 1,2

OpenStudy (anonymous):

and don't forget the plus or minus looks like \[\pm1,\pm2,\pm5,\pm10,\pm\frac{1}{2},\pm\frac{5}{2}\] did i get them all?

OpenStudy (amistre64):

looks about right

OpenStudy (anonymous):

thsats all u do?

OpenStudy (amistre64):

1,2,5,10 +- -------- = 1,2,5,10 1 1,2,5,10 +- -------- = 1/2, 2/2, 5/2, 10/2 2 if a number doubles or more, only count it once 1, 2, 5,10 1/2, 1, 5/2, 5 --------------------- +- 1/2, 1, 2, 5/2, 5, 10

OpenStudy (amistre64):

thats all you do

OpenStudy (anonymous):

Hey Amistre, pls look into my problem too.......